Ammonium nitrate, a common fertilizer, is used as an explosive in fireworks and by terrorists. It was the material used in the tragic explosion at the Oklahoma City federal building in 1995. How many liters of gas at 307°C and 1.00 atm are formed by the explosive decomposition of 80.3 kg of ammonium nitrate to nitrogen, oxygen, and water vapor? Do not answer in scientific notation.

1
Expert's answer
2021-03-10T06:04:58-0500

80.3 x 103 g x mol y mol

2 NH4NO3 = 2 N2 + O2 + 4 H2O

2 x 80 g/mol 2 mol 1 mol


(80.3 x 103) / (2 x 80) = x / 2

x = 1003.75 mol (N2)

(80.3 x 103) / (2 x 80) = x / 1

x = 501.875 mol (O2)

n(gas) = 1003.75 mol + 501.875 mol = 1505.625 mol

pV = nRT

1.00 atm x V = 1505.625 mol x 0.082 L⋅atm⋅K−1⋅mol−1 x (307 + 273) K

V = 71 607.525 L